OpenStudy (anonymous):

How exactly do I solve this? Find the common logarithm of cos81degrees 40'. a. 9.1612 - 10 b. 9.9954 - 10 c. 9.8388 - 10 d. 9.0046 - 10

OpenStudy (lgbasallote):

common log? not common ANTIlog?

OpenStudy (lgbasallote):

then it's just \[\log_{10} (\cos 81^o 40')\]right?

jimthompson5910 (jim_thompson5910):

60 minutes = 1 degree 1 minute = 1/60 degrees 40 minutes = 40/60 degrees 40 minutes = 0.667 degrees So 81degrees 40' = 81.667 degrees roughly

OpenStudy (anonymous):

yikes this really musts be an old old book

OpenStudy (anonymous):

must be the budget cuts

OpenStudy (anonymous):

are you supposed to be looking up these answers in a table in the back of the book? lawda mercy

OpenStudy (anonymous):

would have helped if the answer is close to the options provided

OpenStudy (anonymous):

no? let me check

OpenStudy (anonymous):

Hmmm. I'm trying to match up the answers

OpenStudy (anonymous):

would have helped more if i use 81 instead of 80

OpenStudy (phi):

type log(cos(81.6666667 deg)))+10= in the google search window

OpenStudy (anonymous):

replace 80 by 81 in wolfram

OpenStudy (anonymous):

>.<

OpenStudy (phi):

remember wolfram uses log10 for log base 10.

OpenStudy (anonymous):

9.16

OpenStudy (anonymous):

it uses natural log by default

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

man this is a pain in the neck!

OpenStudy (anonymous):

yeah,the first one

OpenStudy (anonymous):

if +10 is needed in the end, then B is the closest

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

i think it is a)

OpenStudy (anonymous):

or A

OpenStudy (anonymous):

charlie, what's the textbook you're using?

OpenStudy (anonymous):

wolf gave me -0.8388 added 10 got 9.1612

OpenStudy (anonymous):

whoaaa that's not what I got... -1.93....

OpenStudy (anonymous):

i think the text is from maybe 1964 or earlier, before the invention of the hand held calculator

OpenStudy (anonymous):

lol. I remember my professor showing me a table of logs from the 60's

OpenStudy (anonymous):

where's my slide rule?

OpenStudy (anonymous):

LOL I'm taking an online Precalculus/Trig class. THANKS SO MUCH GUYS FOR YOUR HELP. Wish I could give more than one medal

OpenStudy (anonymous):

base of 10 is .0.83

OpenStudy (anonymous):

-0.83

OpenStudy (anonymous):

on line!!!! wtf

OpenStudy (anonymous):

no medal for me please...

OpenStudy (anonymous):

ancient history math with an online course no irony here

OpenStudy (anonymous):

@panlac01 YOU DESERVE ONE I'm sorry :(

OpenStudy (anonymous):

i sent one for you

OpenStudy (anonymous):

yeah. I was taking a hybrid class with face-to-face lecture and then hw's and test were done onine

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

charlie, my textbook is prolly what you need it's exactly the name of the course - pre cal with trig

OpenStudy (anonymous):

Hmm if only I could reach through this laptop screen and steal that book from you haha

OpenStudy (anonymous):

I've sent them to you, but you don't use them....

OpenStudy (anonymous):

That page was difficult to understand!

OpenStudy (anonymous):

http://finedrafts.com/files/Larson%20PreCal%208th/

OpenStudy (anonymous):

O.k. thanks!

OpenStudy (anonymous):

then you probably need intermediate algebra